# Neural network prediction of model parameters for strong lensing samples from Hyper Suprime-Cam Survey

### v2

@@ Abstract
Strong lensing of background galaxies provides important information about the matter distribution around lens galaxies. Traditional modelling of such strong lenses is both time and resource intensive. Fast and automated analysis methods are the need of the hour given large upcoming surveys. In this work, we build and train a simple convolutional neural network with an aim of rapidly predicting model parameters of gravitational lenses. We focus on the inference of the Einstein radius, and ellipticity components of the mass distribution. We train our network on a variety of simulated data with increasing degree of realism and compare its performance on simulated test data in a quantitative manner. We also model 182 gravitational lenses from Subaru HSC survey using YattaLens pipeline to infer their model parameters, which allow a benchmark to compare the predictions of the network. Given all considerations, we conclude that the network trained on simulated samples with lensed sources injected in empty HSC cutouts is the most robust, reproducing Einstein radii with an accuracy of about $10-20$ percent, a bias less than 5 percent, and an outlier fraction of the order of 10 percent. We argue in favour of the subtraction of the lens light before modelling the lens mass distribution. Our comparisons of the inferred parameters of 10 HSC lenses previously modelled in the literature, demonstrate agreement on the Einstein radius. However, the ellipticity components from the network as well as the individual modelling methods, seem to have systematic uncertainties beyond the quoted errors.
@@
pubyear: 2015pagerange: Neural network prediction of model parameters for strong lensing samples from Hyper Suprime-Cam Survey2
@@ 1 Introduction
Lens systems have been searched in ongoing imaging surveys such as the Hyper Suprime-Cam (HSC) survey through campaigns such as the Survey of Gravitationally-lensed Objects in HSC Imaging (SuGOHI). This has resulted in the discovery of hundreds of strong gravitational lens candidates in various categories using different search algorithms (see table 1 in Jaelani et al., 2023). Lens systems have also been searched in the Dark Energy survey via a variety of techniques (e.g., Nord et al., 2020; Rojas et al., 2022). The number of discovered gravitational lens systems is going to increase by an order of magnitude with the next generation ground based imaging surveys such as the Vera Rubin Observatory’s Legacy Survey of Space and Time (LSST). Lens modelling techniques (e.g., Nightingale et al., 2021) are used to estimate the parameters that describe the mass distribution in such individual lens systems. Lens modelling is computationally expensive, often requires attention to individual systems, in addition to development of sophisticated lens modelling codes. Some of the steps that require human attention are identifying lensed features, masking out foreground contaminants, and finding an adequate initial guess for the lens model parameters. Finding more efficient ways to study gravitational lenses is the need of the hour given the large amount of data which will be available imminently.
@@
On the other hand, the era of large scale computing has already begun. Machine Learning (ML) has emerged as a revolutionary tool in order to mine data and identify features or various characteristics of the data (e.g., Moriwaki et al., 2023). The automation implies faster analysis and results. ML techniques such as neural networks are a very effective way in performing tasks like image classifications and parameter estimations. A neural network can be trained to analyse images of gravitationally lensed systems, for purposes like finding out strong lenses (e.g., Jacobs et al., 2017; Rojas et al., 2022) and estimating lens model parameters (e.g., Hezaveh et al., 2017; Pearson et al., 2019; Pearson et al., 2021; Schuldt et al., 2021, 2023a; Schuldt et al., 2023b; Gentile et al., 2023). A neural network once trained, can predict the parameters just by analysing the features in the image such as shape, image separation and brightness. A neural network can statistically infer the parameters of commonly used lensed models, with accuracy comparable to these sophisticated methods but about ten million times faster (Hezaveh et al., 2017). Once we infer the lens parameters statistically, we can use them to probe various dark matter properties, for instance, the dark matter mass fraction (Oguri et al., 2014) as a function of distance away from lensing galaxies and mass distribution of lens systems (e.g., More et al., 2012). ML approach may not lead to accurate measurements on individual lens systems but can provide reasonable estimates for the distribution of lens parameters in a statistical manner.
@@
Hezaveh et al. (2017) demonstrated that they could use Convolutional Neural Network (CNNs) to infer the lens model parameters, in particular, the Einstein radius. Furthermore, Levasseur et al. (2017) have applied a Bayesian framework to determine the uncertainties in the lens parameters. These studies have mainly considered images of strong lenses at high spatial resolution, for instance, with image quality similar to that of the Hubble Space Telescope.
@@
However, a significant fraction of strong lens systems in the near future will be obtained from the images taken by ground based telescopes, where we will have to deal with unique challenges including poorer image quality due to atmospheric seeing and the lower pixel resolution. To address this, we have designed a CNN that enables us to analyse the images of lens systems taken from ground based telescope surveys. A study of this nature has been carried out by Pearson et al. (2019, henceforth, referred to as P19) focussing on simulations for lensed images detectable in ground based Rubin LSST as well as the space based Euclid mission. In our study, we focus on the Subaru Hyper Suprime-Cam survey (HSC) given the existence of data into which simulated lens systems could be injected, a sample of known lenses from the survey, and its similarity in terms of depth and image quality to the Rubin LSST. We train our CNN on a simulated sample of HSC-like lenses and test it on the simulated as well as real grade A and B SuGOHI lenses to infer lens mass model parameters like the Einstein radius, the axis ratio and the position angle of the major axis of the lens mass distribution. Unlike P19, who assumed perfect lens subtraction, we have to process the real SuGOHI lenses before feeding them to the network using a pipeline called YattaLens (YL , Sonnenfeld et al., 2018) in order to subtract the lens light and infer the lens model parameters to obtain the benchmarks for our study.
@@
In Schuldt et al. (2021, henceforth, referred to as S21), they present a simple neural network to infer the lens model parameters of the simulated data, matched to the quality of ground based imaging data from the HSC survey. During the course of this work, Schuldt et al. (2023a, henceforth, referred to as S23a) presented a residual network (ResNet) for the same and in Schuldt et al. (2023b, henceforth, referred to as S23b) they further applied the network on the real SuGOHI lenses. Even though there are a number of similarities between our approaches, there are a number of differences in terms of the choice of the network, the suite and diversity of training samples, their weights and the test samples. In this study, we will also systematically analyse the reasons for the various failure modes of the network and show quantitative estimates of the performance of our network trained on a set of simulations with increasing amount of realism. Similar to S23b, we will also compare our results with the inferences from traditional modelling carried out in Sonnenfeld et al. (2019, henceforth, referred to as S19) and in S23b for a sample of 10 grade A SuGOHI lenses. Given these differences, our study is thus useful to establish the conclusions from P19 and S23b that are likely independent of the precise choice of network as well as the the impact of differences in training samples.
@@ 2.1 Simulated datasets
The simulated lensed arcs used in this work come from the same sample as that used in Jaelani et al. (2023). The reader is referred to Section 3 of Jaelani et al. (2023) for the specific settings, the scaling relations and the distributions used to produce this sample. We note that we have added Poisson noise to the simulated model arcs after convolving with the coadded PSF from the HSC database available at the location of central lens galaxy in each of the g, r and i bands.
We construct numerous training samples, successively, by adding various degrees of realism to our simulated arcs. These training samples were labelled as the following cases: a) Simulated arcs (PSF-convolved and Poisson noise added) without any background noise (henceforth, referred to as the PureSims sample), b) Simulated arcs, as before, but with Gaussian background noise corresponding to the HSC depth (henceforth, referred to as the GauNoise sample), c)Simulated arcs (same as in sample a) added to selected cutouts from HSC with a size of 17 arcsec on the side. The cutouts are selected such that they do not contain any galaxy in the central region of radius of 5 arcsec and brighter than $24.5$ magnitude in the $i-$ band to avoid contamination of the lensed arcs (henceforth, referred to as the HSCempty sample). d) Simulated arcs (same as in sample a) added to the image cutout of the respective HSC galaxy (henceforth, referred to as the LensLight sample). We will describe the purpose behind these different training samples while discussing the results in Section 4.
@@ 2.2 Real datasets
In addition to the sample of simulated lenses reserved for testing, we also make use of a real life test sample. This test sample of real lenses consists of 182 grade A and B SuGOHI lenses. We are also using a sub-sample of 25 grade A lenses. These real images contain a lens galaxy at the center which in most of the cases outshines or contaminates the background lensed sources. It is mentioned in Hezaveh et al. (2017), where they use the high quality images of strong lenses from Hubble Space Telescope that the subtraction of the central lens light helped in improving their results. We are working with ground based data, where the blending is even more prominent. This prompted us to perform lens light subtraction on the SuGOHI lenses before feeding them to the network and we use the YL pipeline for this purpose. The YL pipeline fits the lens light and subtracts it out, identifies the lensed sources and then fits a SIE lens model to predict the parameters. We are using the lens-light subtracted images and further cleaning them by removing foreground objects identified by the YL segmentation results while constructing our test sample of real lenses. For grade A+B and grade A test samples, we use parameters predicted by YL to compare the predictions from our network. Finally, we find 10 SuGOHI lenses in the literature which have also been modelled by others with different algorithms for which we present a comparative analysis. Compared to P19, our network results thus also incorporate uncertainties due to the imperfections in the lens light subtraction.
@@ 3 Network Architecture and Training
We develop a network following the conventional CNN architecture. Our network has seven convolutional blocks with alternate average pooling and two fully connected layers followed by an output layer. Each convolutional block has a convolutional layer followed by batch normalization, Parametric Rectified Linear Unit (PReLU) activation and a dropout layer with 20 per cent dropout rate. We are using mean squared error (MSE) loss function and the Adam optimizer with a learning rate of 0.001 during training. Following the results of P19, we use data from multiple optical wavelengths and input the g, r and i band images of a strong lens system. We use cutouts of $101\times 101$ pixels for each band and expect the network to perform regression and output the lens mass model parameters we are most interested in, namely the Einstein radius, axis ratio and position angle. In the course of our investigations, we have also tried the addition of offsets between the lens galaxy light and lens potential of the magnitude similar to S21, and found not much of a difference in the network predictions for the parameters of our interest. This is likely due to the small magnitudes of these offsets.
We train our model on a sample of 60000 simulated images of lens systems along with their lens model parameters namely Einstein radius, axis ratio and position angle. These 60000 images are obtained from augmenting unique 20000 images of strongly lensed systems. The distributions of the lens model parameters corresponding to these simulated lenses are obtained from the distributions of HSC lenses and it mimics the real distribution of lenses in the Universe. These distributions are naturally imbalanced (see discussion in S21), and we will use explore the use of weights while we train our network. For instance, large Einstein radius systems are expected to be very few in number, which could result in poor training in the corresponding parameter range. At the same time, some of the parameters could be difficult to train than others, like ellipicity in our case. Therefore, we have also explored the use of sample weights (i.e., weights corresponding to each training sample) and class weights (i.e., weights across the different parameters of a training sample). These weights modify the MSE loss during training, such that
@@
where $N$ is the total number of images in the batch, $w_{i}$ is the sample weight for the $i^{\rm{th}}$ image, $\tilde{w}_{j}$ is the class weight for the $j^{\rm{th}}$ parameter, $\theta_{j}$ represents true value of the $j^{\rm{th}}$parameter and $\tilde{\theta}_{j}$ corresponds to its value predicted by the network. We test for multiple values of sample weights and class weights monitoring the variance, the skewness and the kurtosis of the difference between the parameters predicted by our network and the underlying truth for our test samples. Based on these tests, we use sample weights equivalent to the square of the parameter value for the Einstein radius to counterbalance the deficiency of lenses with a higher Einstein radius. We use class weights with values equal to 1, 10 and 5 for the Einstein radius, axis ratio and position angle, respectively.
@@
Figure 2: Performance of our network trained with the PureSims sample (60000 images) on the corresponding test sample of 20000 lenses (upper row) and a test sample of 182 SuGOHI (grade A+B) lenses processed using YL to remove the central lens light and foreground objects (lower row). image cutouts (first column) s 101$\times$101 pixels corresponding to an angular size of $\sim~{}17$ arcsec $\times$ $17$ arcsec. The second, third and fourth columns show the network predictions for Einstein radius ($\theta_{\rm E}$), ellipticity components ($e_{x}$ and $e_{y}$), respectively, as compared to the true values. The range for every parameter is divided into 10 bins (logarithmic for $\theta_{\rm E}$ and linear for $e_{x}$ and $e_{y}$). The red curve represents the median and the dark gray and the light gray regions correspond to the 68 and 95 percentiles, respectively.
@@
Figure 7: Comparing the predictions of our network with true lens light ellipticity (upper row) and true lens mass ellipticity (lower row). In this case, we train our network on a training sample (similar to the LensLight sample) of 40000 lenses, with larger offsets between the lens light and the lens mass parameters ( $\sigma_{\rm q}=0.01$ and $\sigma_{\rm PA}=30$°) and test it on the corresponding test sample consisting of 1781 lenses.
Figure 8: A sample of colour-composite images of 10 SuGOHI lenses commonly modelled in S19, S23b, YL and our work. We have processed these lenses using YL in order to remove the central lens light and foreground objects. Columns from left to right : original image, lens-subtracted image, arc and image segmentation map (green = arcs, white = modelled foreground objects, red = masked out foreground objects), best-fitting lens model of the system (including modelled foreground objects), best model of the lensed background galaxy alone (source only model), residuals between the data and the best-fitting lens model.
Figure 9: Comparison of the the lens model parameters, namely, Einstein radius ($\theta_{\rm E}$) and ellipticity components ($e_{x}$ and $e_{y}$ ), predicted by our network (this work) with YL , S19 and S23b results. The error bars on our network predictions are obtained as described in the Appendix  B. We convert the results of YL , S19 and S23b to our conventions before comparing (see Appendix A).

### v3

@@ Abstract
Strong lensing of background galaxies provides important information about the matter distribution around lens galaxies. Traditional modelling of such strong lenses is both time and resource intensive. Fast and automated analysis methods are the need of the hour given large upcoming surveys. In this work, we build and train a simple convolutional neural network with an aim of rapidly predicting model parameters of gravitational lenses. We focus on the inference of the Einstein radius, and ellipticity components of the mass distribution. We train our network on a variety of simulated data with increasing degree of realism and compare its performance on simulated test data in a quantitative manner. We also model 182 gravitational lenses from the HSC survey using YattaLens pipeline to infer their model parameters, which allow a benchmark to compare the predictions of the network. Given all considerations, we conclude that the network trained on simulated samples with lensed sources injected in empty HSC cutouts is the most robust, reproducing Einstein radii with an accuracy of about $10-20$ percent, a bias less than 5 percent, and an outlier fraction of the order of 10 percent. We argue in favour of the subtraction of the lens light before modelling the lens mass distribution. Our comparisons of the inferred parameters of 10 HSC lenses previously modelled in the literature, demonstrate agreement on the Einstein radius. However, the ellipticity components from the network as well as the individual modelling methods, seem to have systematic uncertainties beyond the quoted errors.
@@
pubyear: 2015pagerange: Neural network prediction of model parameters for strong lensing samples from Hyper Suprime-Cam Survey13
@@ 1 Introduction
Lens systems have been searched in ongoing imaging surveys such as the Subaru Hyper Suprime-Cam (HSC) survey through campaigns such as the Survey of Gravitationally-lensed Objects in HSC Imaging (SuGOHI). This has resulted in the discovery of hundreds of strong gravitational lens candidates in various categories using different search algorithms (see table 1 in Jaelani et al., 2024). Lens systems have also been searched in surveys including DES (e.g., O’Donnell et al., 2022), SLACS (e.g., Bolton et al., 2006), SL2S (e.g., Sonnenfeld et al., 2013) and more recently, Euclid (e.g., Nagam et al., 2025).
@@
The number of discovered gravitational lens systems is going to increase by an order of magnitude with the next generation ground based imaging surveys (e.g., Collett, 2015) such as the Vera Rubin Observatory’s Legacy Survey of Space and Time (LSST). Forward lens modelling techniques (e.g., Nightingale et al., 2021; Birrer & Amara, 2018; Suyu & Halkola, 2010) are used to estimate the parameters that describe the mass distribution in such individual lens systems. Lens modelling is computationally expensive, often requires attention to individual systems, in addition to development of sophisticated lens modelling codes. Some of the steps that require human attention are identifying lensed features, masking out foreground contaminants, and finding an adequate initial guess for the lens model parameters. Finding more efficient ways to study gravitational lenses is the need of the hour given the large amount of data which will be available imminently.
@@
On the other hand, the era of large scale computing has already begun. Machine Learning (ML) has emerged as a revolutionary tool in order to mine data and identify features or various characteristics of the data (e.g., Moriwaki et al., 2023). The automation implies faster analysis and results. ML techniques such as neural networks are a very effective way in performing tasks like image classifications and parameter estimations. A neural network can be trained to analyse images of gravitationally lensed systems, for purposes like finding out strong lenses (e.g., Jacobs et al., 2017; Rojas et al., 2022) and estimating lens model parameters (e.g., Hezaveh et al., 2017; Pearson et al., 2019; Pearson et al., 2021; Schuldt et al., 2021, 2023a; Schuldt et al., 2023b; Gentile et al., 2023). A neural network once trained, can predict the parameters just by analysing the features in the image such as shape, image separation and brightness. A neural network can statistically infer the parameters of commonly used lensed models, with accuracy comparable to these sophisticated methods but about ten million times faster (Hezaveh et al., 2017). Once we infer the lens parameters statistically, we can use them to probe various dark matter properties, for instance, the dark matter mass fraction (Oguri et al., 2014) as a function of distance away from lensing galaxies and mass distribution of lens systems (e.g., More et al., 2012).
@@
Hezaveh et al. (2017) demonstrated that they could use Convolutional Neural Network (CNNs) to infer the lens model parameters, in particular, the Einstein radius. Furthermore, Perreault Levasseur et al. (2017) have applied a Bayesian framework to determine the uncertainties in the lens parameters. These studies have mainly considered images of strong lenses at high spatial resolution, for instance, with image quality similar to that of the Hubble Space Telescope.
@@
However, a significant fraction of strong lens systems in the near future will be obtained from the images taken by ground based telescopes, where we will have to deal with unique challenges including poorer image quality due to atmospheric seeing and the lower pixel resolution. To address this, we have designed a CNN that enables us to analyse the images of lens systems taken from ground based telescope surveys. A study of this nature has been carried out by Pearson et al. (2019, henceforth, referred to as P19) focussing on simulations for lensed images detectable in ground based LSST as well as the space based Euclid mission. In our study, we focus on the HSC survey given the existence of data into which simulated lens systems could be injected, a sample of known lenses from the survey, and its similarity in terms of depth and image quality to the LSST. We train our CNN on a simulated sample of HSC-like lenses and test it on the simulated as well as real grade A and B SuGOHI lenses to infer lens mass model parameters like the Einstein radius, the axis ratio and the position angle of the major axis of the lens mass distribution. Unlike P19, who assumed perfect lens subtraction, we have to process the real SuGOHI lenses before feeding them to the network using a pipeline called YattaLens (YL , Sonnenfeld et al., 2018) in order to subtract the lens light and infer the lens model parameters to obtain the benchmarks for our study.
In Schuldt et al. (2021, henceforth, referred to as S21), they present a simple neural network to infer the lens model parameters of the simulated data, matched to the quality of ground based imaging data from the HSC survey. During the course of this work, Schuldt et al. (2023a, henceforth, referred to as S23a) presented a residual network (ResNet) for the same and in Schuldt et al. (2023b, henceforth, referred to as S23b) they further applied the network on the real SuGOHI lenses.
@@
Even though there are a number of similarities between our approaches, there are a number of differences in terms of the choice of the network, the suite and diversity of training samples, their weights and the test samples. In this study, we will also systematically analyse the reasons for the various failure modes of the network and show quantitative estimates of the performance of our network trained on a set of simulations with increasing amount of realism. Similar to S23b, we will also compare our results with the inferences from traditional modelling carried out in Sonnenfeld et al. (2019, henceforth, referred to as S19) and in S23b for a sample of 10 grade A SuGOHI lenses. Given these differences, our study is thus useful to establish the conclusions from P19 and S23b that are likely independent of the precise choice of network as well as the the impact of differences in training samples.
@@ 2.1 Simulated datasets
The simulated lensed arcs used in this work come from the same sample as that used in Jaelani et al. (2024). The reader is referred to Section 3 of Jaelani et al. (2024) for the specific settings, the scaling relations and the distributions used to produce this sample. In Fig. 2, we present the corresponding distributions of the lens mass model parameters under consideration. We note that we have added Poisson noise to the simulated model arcs after convolving with the coadded PSF from the HSC database available at the location of central lens galaxy in each of the g, r and i bands.
We construct numerous training samples, successively, by adding various degrees of realism to our simulated arcs. These training samples were labelled as the following cases: a) Simulated arcs (PSF-convolved and Poisson noise added) without any background noise (henceforth, referred to as the PureSims sample), b) Simulated arcs, as before, but with Gaussian background noise corresponding to the HSC depth (henceforth, referred to as the GauNoise sample), c)Simulated arcs (same as in sample a) added to selected cutouts from HSC with a size of 17 arcsec on the side. The cutouts are selected such that they do not contain any galaxy in the central region of radius of 5 arcsec and brighter than $24.5$ magnitude in the $i-$ band to avoid contamination of the lensed arcs. The images in this sample resemble real lenses after the ideal subtraction of the lens light (henceforth, referred to as the HSCempty sample). d) Simulated arcs (same as in sample a) added to the image cutout of the respective HSC galaxy (henceforth, referred to as the LensLight sample). We would like to highlight that all of the above samples, except for the LensLight sample, contain images without the lens light. We will describe the purpose behind these different training samples while discussing the results in Section 4.
@@ 2.2 Real datasets
In addition to the sample of simulated lenses reserved for testing, we also make use of a real life test sample. This test sample of real lenses consists of 182 grade A and B SuGOHI lenses. We are also using a sub-sample of 25 grade A lenses. These real images contain a lens galaxy at the center which in most of the cases outshines or contaminates the background lensed sources. It is mentioned in Hezaveh et al. (2017), where they use the high quality images of strong lenses from Hubble Space Telescope that the subtraction of the central lens light helped in improving their results. We are working with ground based data, where the blending is even more prominent. This prompted us to perform lens light subtraction on the SuGOHI lenses before feeding them to the network and we use the YL pipeline for this purpose. The YL pipeline fits the lens light and subtracts it out, identifies the lensed sources and then fits a SIE lens model to predict the parameters. We use lens-light subtracted images and further clean them by removing foreground objects identified by the YL segmentation results while constructing our test sample of real lenses. For grade A+B and grade A test samples, we use parameters predicted by YL to compare the predictions from our network. Finally, we find 10 SuGOHI lenses in the literature which have also been modelled by others with different algorithms for which we present a comparative analysis.
@@ 3 Network Architecture and Training
We develop a network following the conventional CNN architecture. Our network has seven convolutional blocks with alternate average pooling and two fully connected layers followed by an output layer. Each convolutional block has a convolutional layer followed by batch normalization, Parametric Rectified Linear Unit (PReLU) activation and a dropout layer with 20 per cent dropout rate. We use a mean-squared error (MSE) loss function and the Adam optimizer with a learning rate of 0.001 during training. Following the results of P19, we use data from multiple optical wavelengths and input the g, r and i band images of a strong lens system. We use cutouts of $101\times 101$ pixels for each band and expect the network to perform regression and output the lens mass model parameters we are most interested in, namely the Einstein radius, axis ratio and position angle. In the course of our investigations, we have also tried the addition of offsets between the lens galaxy light and lens potential of the magnitude similar to S21, and found not much of a difference in the network predictions for the parameters of our interest. This is likely due to the small magnitudes of these offsets.
We train our model on a sample of 60000 simulated images of lens systems along with their lens model parameters namely Einstein radius, axis ratio and position angle. These 60000 images are obtained from augmenting 20000 unique images of strongly lensed systems by applying rotation. The distributions of the lens model parameters corresponding to these simulated lenses are obtained from the distributions of HSC lenses and mimic the real distribution of lenses in the Universe. These distributions are naturally imbalanced for certain parameters (see discussion in S21), and we will use explore the use of weights while we train our network. For instance, large Einstein radius systems are expected to be very few in number, which could result in poor training in the corresponding parameter range. At the same time, some of the parameters, such as ellipticity, are more difficult for a network to learn. Therefore, we have also explored the use of sample weights (i.e., weights corresponding to each training sample) and class weights (i.e., weights across the different parameters of a training sample). These weights modify the MSE loss during training, such that
@@
where $N$ is the total number of images in the batch, $w_{i}$ is the sample weight for the $i^{\rm{th}}$ image, $\tilde{w}_{j}$ is the class weight for the $j^{\rm{th}}$ parameter, $\theta_{j}$ represents true value of the $j^{\rm{th}}$parameter and $\tilde{\theta}_{j}$ corresponds to its value predicted by the network. We test for multiple values of sample weights and class weights monitoring the variance, the skewness and the kurtosis of the difference between the parameters predicted by our network and the underlying truth for our test samples. Based on these tests, we use sample weights equivalent to the square of the parameter value for the Einstein radius to counterbalance the lack of lenses with a higher Einstein radius (see Fig. 2). We use class weights with values equal to 1, 10 and 5 for the Einstein radius, axis ratio and position angle, respectively.
@@
Figure 2: The blue histograms show the normalized distributions of the parameters $\theta_{\rm E}$, $e_{x}$ and $e_{y}$ for the training sample. For $\theta_{\rm E}$, the orange histogram represents the distribution obtained after applying sample weights equivalent to $\theta_{\rm E}^{2}$, highlighting the effect of weighting to counterbalance the lack of lenses with higher Einstein radii.
Figure 3: Performance of our network trained with the PureSims sample (60000 images) on the corresponding test sample of 20000 lenses (upper row) and a test sample of 182 SuGOHI (grade A+B) lenses processed using YL to remove the central lens light and foreground objects (lower row). The size of the image cutouts (first column) is 101$\times$101 pixels corresponding to an angular size of $\sim~{}17$ arcsec $\times$ $17$ arcsec. The second, third and fourth columns show the network predictions for Einstein radius ($\theta_{\rm E}$), ellipticity components ($e_{x}$ and $e_{y}$), respectively, as compared to the true values. The range for every parameter is divided into 10 bins (logarithmic for $\theta_{\rm E}$ and linear for $e_{x}$ and $e_{y}$). The red curve represents the median and the dark gray and the light gray regions correspond to the 68 and 95 percentiles, respectively.
@@
Figure 8: Comparing the predictions of our network with true lens light ellipticity (upper row) and true lens mass ellipticity (lower row). In this case, we train our network on a training sample (similar to the LensLight sample) of 40000 lenses, with larger offsets between the lens light and the lens mass parameters ( $\sigma_{\rm q}=0.1$ and $\sigma_{\rm PA}=30$°) and test it on the corresponding test sample consisting of 1781 lenses.
Figure 9: A sample of colour-composite images of 10 SuGOHI lenses commonly modelled in S19, S23b, YL and our work. We have processed these lenses using YL in order to remove the central lens light and foreground objects. Columns from left to right : original image, lens light-subtracted image, arc and image segmentation map (green = arcs, white = modelled foreground objects, red = masked out foreground objects), best-fitting lens model of the system (including modelled foreground objects), best model of the lensed background galaxy alone (source only model), residuals between the data and the best-fitting lens model.
Figure 10: Comparison of the the lens model parameters, namely, Einstein radius ($\theta_{\rm E}$) and ellipticity components ($e_{x}$ and $e_{y}$ ), predicted by our network (this work) with YL , S19 and S23b results. The error bars on our network predictions are obtained as described in the Appendix  B. We convert the results of YL , S19 and S23b to our conventions before comparing (see Appendix A). In S23b, asymmetric uncertainties are reported for $\theta_{\rm E}$, $q$, and PA. We adopt the conservative bounds and propagate these errors assuming Gaussian distributions and treating the three parameters as independent. In contrast, S19 assumes constant uncertainties for each parameter ($\theta_{\rm E}$: 0.02 arcsec, $q$: 0.01, PA: 0.8°), which we omit from the plot for clarity.